Halliford Corporation expects to have earnings this coming year of $ 2.784$2.784 per share. Halliford plans to retain all of its earnings for the next two years.​ Then, for the subsequent two​ years, the firm will retain 53 %53% of its earnings. It will retain 21 %21% of its earnings from that point onward. Each​ year, retained earnings will be invested in new projects with an expected return of 22.1 %22.1% per year. Any earnings that are not retained will be paid out as dividends. Assume​ Halliford's share count remains constant and all earnings growth comes from the investment of retained earnings. If​ Halliford's equity cost of capital is 11.4 %11.4%​, what price would you estimate for Halliford​ stock?
Answer #1

The growth rate = Retention ratio * Expected return

Growth rate for first two years = 1*0.221 = 0.221

Growth rate for years 3 and 4 = 0.53*0.221 = 0.11713

Growth rate from year 5 onwards = 0.21*0.221 = 0.04641

The calculation of dividends is as shown in the table below:

Year 0 1 2 3 4 5 6
EPS Growth rate 0.221 0.221 0.11713 0.11713 0.04641
EPS 2.784 3.399264 4.150501 4.63665 5.17974 5.420132
Retention Ratio 100% 100% 53% 53% 21% 21%
Dividend Payout 0.00% 0.00% 47.00% 47.00% 79.00% 79.00%
Actual dividend 0 0 1.950736 2.179225 4.091995 4.281904

Price in year 5 = D6/(r-g) = 4.81904/(0.114-0.005) = 39.28353

Current Price (P0) = 1.950736/1.114^3 + 2.179225/1.114^4 + 4.091995/1.114^5 + 39.28353/1.114^5

Current Price (P0) =$28.11 (Rounded to 2 decimals)